### v2

@@ 4.3 Comparison with State-of-the-Arts
Person Re-Identification. To further prove the effectiveness of our method on other related object ReID tasks, we also compare our DMCL with existing methods in Table 2. we can easily observe that our method can achieve the state-of-the-art performance on Market-1501, DukeMTMC-reID and CUHK03 datasets with considerable advantages respectively. For example, our DMCL method can achieve a mAP/Rank-1 performance of 92.1% and 96.3% respectively on Market-1501 dataset, leading +1.0% and +0.6% improvement of mAP and Rank-1 accuracy when compared to the second best method NFormer (Wang et al, 2022) and MGN (Wang et al, 2018). Besides, our DMCL method can also obtain the improvement of +1.2% and +1.4% in terms of mAP and Rank-1 accuracy on CUHK03 dataset when compared to the VT-ReID (Xiang et al, 2024b). These experiments strongly demonstrates the priority of our proposed deep multimodal collaborative learning framework.
Table 2: Performance comparison with other methods on Person ReID datasets. Bold indicates the best and underline the second best.
@@
4.4 Ablation Studies
Effectiveness of deep multimodal collaborative learning: Firstly, from the quantitative aspect, we evaluate the effectiveness of deep multimodal collaborative learning DMCL framework. As illustrated in Table 3, when adopting our deep multimodal collaborative learning framework, results show that mAP accuracy improves significantly from 25.9% to 46.4% on Colo-Pair dataset with visual-text induction. Additionally, similar improvements can also be easily observed in terms of Rank-1, Rank-5 evaluation metrics, leading to +24.3% and +9.1% respectively. These results prove that multimodal collaborative learning paradigm has a direct impact on downstream polyp ReID task.
@@ 4.4 Ablation Studies
DMCL w/ text $\checkmark$ $\times$ 28.5 18.6 38.8
DMCL w/ image $\times$ $\checkmark$ 36.6 41.8 46.2
@@
Figure 3: Qualitative visualization of ranking results of our proposed approach DMCL with collaborative training mechanism on Colo-ReID dataset.
Secondly, from the qualitative aspect, we also give some qualitative results of our proposed deep multimodal collaborative learning framework named DMCL. Fig. 3 provides some qualitative visualization of ranking results of our proposed approach DMCL with collaborative training mechanism on Colo-ReID dataset. To be more specific, we can obviously observe that our model attends to relevant image regions or discriminative parts for making polyp predictions, indicating that DMCL can greatly help the model learn more global context information and meaningful visual features with better semantic understanding, which significantly makes our collaborative training method DMCL model more robust to perturbations.
Effectiveness of dynamic multimodal training strategy: In this section, we proceed to evaluate the effectiveness of dynamic multimodal training strategy by testing whether text modality or image modality matters. According to Table 3, our dynamic multimodal training strategy DMCL with text representation (DMCL w/ text) can lead to a significant improvement in Rank-1 of +5.6% on Colo-Pair dataset when compared with baseline setting. Furthermore, when adopting image representation, our method (DMCL w/ image) can obtain a remarkable performance of 36.6% in terms of mAP accuracy, leadig a significant improvement of +8.1% when compared to DMCL w/ text. The effectiveness of the dynamic multimodal training strategy can be largely attributed to that it enhances the discrimination capability of other collaborative networks during multimodal representation learning, which is vital for polyp domain adaptation in general domain where the target supervision is not available.
@@
Visualization of Feature Response: To further explain why our deep multimodal collaborative learning strategy works, we perform in-depth analysis of feature response in DMCL method, and also show some qualitative examples of EigenGradCAM (Muhammad and Yeasin, 2020) visualizations in Fig.4. In fact, the Grad-CAM is a package with some methods for Explainable AI for computer vision, which provides attributions to both the inputs and the neurons of intermediate layers, so as to make CNN-based models more transparent by producing visual explanations. Specifically, EigenGradCAM decodes the importance of each feature map to a specific class by analyzing the gradients within the last convolutional layer of a CNN, generating a heatmap that highlights the image regions that contribute most significantly to the prediction result. This heatmap visually represents which parts of the image are most influential in the model’s final decision. As illustrated in Fig.4, compared with CNN-based training method without text information, we observe that our model attends to relevant image regions or discriminative parts for making decison, indicating that our DMCL model can effectively explore more global context information and meaningful visual features with better semantic understanding, which significantly make our model more robust to perturbations.
@@ 5 Conclusion
This study further investigates the possibility of applying multimodal collaborative learning for polyp retrieval task with visual-text dataset, and then propose a simple but effective multimodal representation learning network named DMCL to improve the performance of polyp Re-ID event. To further enhance the robustness of DMCL model, a dynamic multimodal feature fusion strategy is introduced to leverage the optimized multimodal representations for multimodal fusion via end-to-end training. Comprehensive experiments on standard benchmarks also demonstrate that our method can achieve the state-of-the-art performance in polyp Re-ID task and other image retriveal task. In the future, we will explore the interpretability of this method and apply it to other related computer vision tasks, e.g. polyp detection and segmentation.
@@ Acknowledgments
This work was partially supported by the National Natural Science Foundation of China under Grant No.62301315, Startup Fund for Young Faculty at SJTU (SFYF at SJTU) under Grant No.23X010501967 and Shanghai Municipal Health Commission Health Industry Clinical Research Special Project under Grant No.202340010. The authors would like to thank the anonymous reviewers for their valuable suggestions and constructive criticisms.
@@
    This work was partially supported by the National Natural Science Foundation of China under Grant No.62301315, Startup Fund for Young Faculty at SJTU (SFYF at SJTU) under Grant No.23X010501967 and Shanghai Municipal Health Commission Health Industry Clinical Research Special Project under Grant No.202340010.
@@
    Suncheng Xiang and Jiacheng Ruan contributed conception and design of the study, as well as the experimental process and interpreted model results. Suncheng Xiang obtained funding for the project and provided clinical guidance. Suncheng Xiang drafted the manuscript. All authors contributed to manuscript revision, read and approved the submitted version.

### v3

@@ 4.3 Comparison with State-of-the-Arts
Person Re-Identification. To further prove the effectiveness of our method on other related object ReID tasks, we also compare our DMCL with existing methods in Table 2. we can easily observe that our method can achieve the state-of-the-art performance on Market-1501, DukeMTMC-reID and CUHK03 datasets with considerable advantages respectively. For example, our DMCL method can achieve a mAP/Rank-1 performance of 92.1% and 96.3% respectively on Market-1501 dataset, leading +1.0% and +0.6% improvement of mAP and Rank-1 accuracy when compared to the second best method NFormer (Wang et al, 2022) and MGN (Wang et al, 2018). In addition, our DMCL method can also obtain the improvement of +1.2% and +1.4% in terms of mAP and Rank-1 accuracy on CUHK03 dataset when compared to the VT-ReID (Xiang et al, 2024). These experiments strongly demonstrate the priority of our proposed deep multimodal collaborative learning framework.
4.4 Ablation Studies
Effectiveness of deep multimodal collaborative learning: Firstly, from the quantitative aspect, we evaluate the effectiveness of deep multimodal collaborative learning DMCL framework. As illustrated in Table 3, when adopting our deep multimodal collaborative learning framework, results show that mAP accuracy improves significantly from 25.9% to 46.4% on Colo-Pair dataset with visual-text induction. Additionally, similar improvements can also be easily observed in terms of Rank-1, Rank-5 evaluation metrics, leading to +36.8% and +20.8% respectively. These results prove that multimodal collaborative learning paradigm has a direct impact on downstream polyp ReID task.
Table 2: Performance comparison with state-of-the-art methods on Person ReID datasets. Bold indicates the best and underline the second best.
@@ 4.4 Ablation Studies
DMCL w/o image $\checkmark$ $\times$ 28.5 18.6 38.8
DMCL w/o text $\times$ $\checkmark$ 36.6 41.8 46.2
@@
Figure 4: The t-SNE visualization of our method with different pre-training settings (e.g. (a) DMCL w/o Text and (b) DMCL) on Colo-Pair dataset. Note that points of the same color represent the same class.
Figure 5: Qualitative visualization of ranking results of our proposed approach DMCL with collaborative training mechanism on Colo-Pair dataset. For each group, the first column shows the query image, while the second and third columns display the inferred results retrieved by our DMCL model. What’s more, we can obviously observe that our method shows a great robustness regardness of the deformation or illumination variation of these captured polyps.
Secondly, from the qualitative aspect, we also give some qualitative results of our proposed deep multimodal collaborative learning framework on multimodal polyp scenarios. For example, according to the t-SNE visualization in Fig. 4, it can be obviously observed that our method with multimodal collaborative learning strategy can aggregate the features of the same polyp more tightly, which can significantly enhance the model’s robustness and generalziation ability. Furthermore, Fig. 5 also provides some qualitative visualization of the ranking results of proposed approach DMCL with collaborative training mechanism on Colo-Pair dataset. To be more specific, we can obviously observe that our model attends to relevant image regions or discriminative parts for making polyp predictions, indicating that DMCL can greatly help the model learn more global context information and meaningful visual features with better multimodal understanding, which provides the possibility to discover deeper layer of neural network and more representative features from images automatically, and significantly makes our collaborative training method DMCL model more adaptable in general polyp scenarios.
Effectiveness of dynamic multimodal training strategy: In this section, we proceed to evaluate the effectiveness of dynamic multimodal training strategy by testing whether text modality or image modality matters. According to Table 3, our dynamic multimodal training strategy DMCL with text representation (DMCL w/o image) can lead to a significant improvement in Rank-1 of +1.1% on Colo-Pair dataset when compared with baseline setting. Furthermore, when adopting image representation, our method (DMCL w/o text) can obtain a remarkable performance of 36.6% in terms of mAP accuracy, leading to a significant improvement of +8.1% when compared to DMCL w/o image. The effectiveness of the dynamic multimodal training strategy can be largely attributed to that it enhances the discrimination capability of collaborative networks during multimodal representation learning, which is vital for polyp re-identification in general domain where the target supervision is not available.
@@
Visualization of Feature Response: To further explain why our deep multimodal collaborative learning strategy works, we perform in-depth analysis of feature response in DMCL method, and also show some qualitative examples of EigenGradCAM (Muhammad and Yeasin, 2020) visualizations in Fig. 6. In fact, the Grad-CAM is a package with some methods for Explainable AI for computer vision, which provides attributions to both the inputs and the neurons of intermediate layers, so as to make CNN-based models more transparent by producing visual explanations. Specifically, EigenGradCAM decodes the importance of each feature map to a specific class by analyzing the gradients within the last convolutional layer of a CNN, generating a heatmap that highlights the image regions that contribute most significantly to the prediction result. This heatmap visually represents which parts of the image are most influential in the model’s final decision. As illustrated in Fig. 6, compared to the CNN-based training method without text information, we observe that our model attends to relevant image regions or discriminative parts for making decisions, indicating that our DMCL model can effectively explore more global context information and meaningful visual features with better semantic understanding, which significantly make our model more robust to perturbations, such as light-colored areas, disturbance due to its contextual discriminability in visual representation.
@@ 5 Conclusion
This study further investigates the possibility of applying multimodal collaborative learning for the polyp retrieval task with visual-text dataset, and then proposes a simple but effective multimodal representation learning network named DMCL to improve the performance of polyp Re-ID event. To further enhance the robustness of DMCL model, a dynamic multimodal feature fusion strategy is introduced to leverage the optimized multimodal representations for multimodal fusion via end-to-end training. Comprehensive experiments on standard benchmarks also demonstrate that our method can achieve the state-of-the-art performance in polyp Re-ID task and other image retrieval task. In the future, we will explore the interpretability of this method and apply it to other related computer vision tasks, e.g. polyp detection and segmentation.
@@ Acknowledgments
This work was partially supported by the National Natural Science Foundation of China under Grant No.62301315, Startup Fund for Young Faculty at SJTU (SFYF at SJTU) under Grant No.23X010501967, Shanghai Municipal Health Commission Health Industry Clinical Research Special Project under Grant No.202340010 and 2025 Key Research Initiatives of Yunnan Erhai Lake National Ecosystem Field Observation Station under Grant No.2025ZD03. The authors would like to thank the anonymous reviewers for their valuable suggestions and constructive criticisms.
@@
    This work was partially supported by the National Natural Science Foundation of China under Grant No.62301315, Startup Fund for Young Faculty at SJTU (SFYF at SJTU) under Grant No.23X010501967, Shanghai Municipal Health Commission Health Industry Clinical Research Special Project under Grant No.202340010 and 2025 Key Research Initiatives of Yunnan Erhai Lake National Ecosystem Field Observation Station under Grant No.2025ZD03.
@@
    Suncheng Xiang and Jiale Guan contributed conception and design of the study, as well as the experimental process and interpreted model results. Suncheng Xiang obtained funding for the project and provided clinical guidance. Suncheng Xiang drafted the manuscript. All authors contributed to manuscript revision, read and approved the submitted version.